### v2

@@ Abstract
Understanding bidding behavior in multi-unit auctions remains an ongoing challenge for researchers. Despite their widespread use, theoretical insights into the bidding behavior, revenue ranking, and efficiency of commonly used multi-unit auctions are limited. This paper utilizes artificial intelligence, specifically reinforcement learning, as a model free learning approach to simulate bidding in three prominent multi-unit auctions employed in practice. We introduce six algorithms that are suitable for learning and bidding in multi-unit auctions and compare them using an illustrative example. This paper underscores the significance of using artificial intelligence in auction design, particularly in enhancing the design of multi-unit auctions.
@@ 1 Introduction
Multi-unit auctions play a fundamental role in allocating goods and services across diverse markets, including treasury bills, emission permits, spectrum licenses, and electricity (Khezr and Cumpston, 2022; Bichler and Goeree, 2017). These auctions are often characterized by asymmetric information and, due to the existence of multiple units, result in a complex bidding environment. Although multi-unit auctions are commonly employed in practice, there is a lack of consensus in the literature regarding their performance in terms of revenue and efficiency. Moreover, economic theory often lacks clear guidance on the outcomes of many multi-unit auctions (Krishna, 2009).
A range of disciplines, including economics and computer science, have employed various methodologies to investigate multi-unit auctions. Auction theory, data analytics, experimental approaches, and simulations represent some of the predominant methods used in these studies (Ausubel et al., 2014; Roughgarden et al., 2017; Morgenstern and Roughgarden, 2016; Hailu and Thoyer, 2007). Each method offers distinct advantages but also comes with limitations. A common shortfall across these approaches is their inability to comparatively assess the performance of different multi-unit auction variations under similar conditions. Additionally, traditional simulation methods often lack clear guidance as they require an existing model or theoretical prediction to feed into the simulation. In this article, we use artificial intelligence, specifically introducing several reinforcement learning (RL) algorithms, to simulate bidding behavior in various multi-unit auctions.
@@
Our approach leverages artificial intelligence as a complementary tool to previous methodologies, aiming to emulate the human thought process and decision-making in complex bidding scenarios within multi-unit auctions. Using a standard private value model, we assume that buyers’ valuations for each unit up for auction are privately known. Our investigation focuses on three sealed-bid multi-unit auctions: the Discriminatory Price (DP), the Generalized Second-Price (GSP), and the Uniform-Price (UP) auctions. All three of these auctions have been employed extensively in real-world markets to sell various goods and services.
We first provide a metric for learning and demonstrate through an example, in which each bidder demands two units, that almost all RL algorithms perform very well in terms of learning to bid. Although they vary in terms of speed and learning patterns, some top-performing algorithms surprisingly converge to very similar strategies in the final episodes of learning. One of the six algorithms, namely Proximal Policy Optimization, not only performs the best by earning the highest payoffs in all cases but also is the most stable in terms of learning. We then compare the revenue and efficiency of the three auctions across three different scenarios.
@@
Since the seminal work by Vickrey (Vickrey, 1961), it is known that bidding behavior in multi-unit auctions does not necessarily follow those in single unit auctions. Despite this, most of conventionally used multi-unit auctions are extensions of a standard single-unit format. For instance, the generalized second-price auction, as appears from its name, is a format that extend the second-price rule to a multi-unit case. There are several studies that show these multi-unit extensions do not necessarily have the ideal properties of the single unit format (e.g. Back and Zender, 1993; Edelman et al., 2007; Ausubel et al., 2014). Additionally, due to multiple equilbria and lack of closed form solutions for the bidding functions, theoretical studies have no clear conclusion about the revenue and efficiency ranking of commonly used multi-unit auctions (Krishna, 2009; Ausubel et al., 2014).
Recently, the use of RL approaches, particularly Q-learning algorithms, has become more prevalent in auction theory literature. For example, Banchio and Skrzypacz (2022) employ a simple Q-learning algorithm to simulate bidding behavior in two different single-unit auction formats. Khezr et al. (2024) is another study that uses Q-learning to simulate a more complex type of auction, namely the knapsack auction. While various other RL approaches could potentially be used to simulate bidding in auctions, current studies are mainly limited to simple Q-learning algorithms. Therefore, our main aim is to explore other, more advanced algorithms that could effectively simulate such complex bidding environments.
@@ 2 Model
The seller uses a sealed-bid multi-unit auction to allocate the objects. Each bidder submits a vector of bids $\mathbf{b}^{i}$ for all the units they demand. The auctioneer ranks the bids from the highest to the lowest, and allocate object to the $K$ highest submitted bids. We explore three types of payment rules as follows.
@@
In this auction, every buyer $i$ submits up to $k$ bids, represented by $\mathbf{b}^{1}=(b_{1}^{i},b_{2}^{i},...,b_{k}^{i})$. The auctioneer arranges all the $n\times k$ bids from the highest to the lowest. The top $k$ bids each win a unit and pay the amount they bid for that particular unit. For example, if the first bid from bidder three, $b_{1}^{3}$, holds the top rank, she secures the first unit and pays $b_{1}^{3}$ for it.
@@
Much like the DP auction, each buyer $i$ submits up to $k$ bids. The top $k$ bids each win a unit.However, in this case, winners pay the amount of the bid that is ranked immediately below their own (excluding their own subsequent bids). For instance, if bidder three’s first bid, $b_{1}^{3}$, holds the highest rank, she secures the first unit but pays the amount of the second-highest bid.
@@
In this auction every winning bidder pays a uniform price for each unit equal to the highest losing bid, that is the $K+1$th highest bid.
@@ 3 Reinforcement Learning approaches
We use Reinforcement Learning (RL) methods to model bidder behavior in multi-unit auctions with multiple bidders. These algorithms are inspired by behavioral psychology, where an agent learns to achieve a goal by receiving feedback through rewards or penalties. An agent, in the context of this work, is a simulated human bidder operating within an environment, an auction, in which they interact with other agents via bidding for available items. The environment encompasses everything the agent interacts with and can change over time, either due to the agent’s actions or independently. Interactions between an agent and the environment are defined by the terms: state, action, and reward.
A state is a complete description of the agent’s situation within the environment. An observation, which might be a partial view of the state, provides the agent with information about the current situation. Actions are the set of all possible decisions the agent can make. A reward is a feedback signal indicating the immediate effect of the agent’s previous action. The agent’s objective is to maximize the cumulative reward, or return, over time. This objective and the associated decision-making strategy are encapsulated in an agent’s policy, a mapping of states (or observations) to actions.
@@ 3.1 Environment setup and evaluation
The reward function, therefore, returns values that can be either positive or negative, rewarding positive payoffs and punishing overbidding more severely than unsuccessful bidding. Specifically, when $p>0$, the agent receives a positive reward as determined by Equation 1. However, when $p<0$, it indicates that the agent’s payment—whether it is their own bid (in DP) or some other lower bid (in GSP and UP)—exceeds their value, thus the numerator of the reward function in Equation 1 becomes negative. In this case, the agent receives a penalty in the form of a negative reward. Both reward and punishment are proportional to the value, which means that for higher values, any given positive or negative payoff is less rewarded or punished, respectively. For instance, if an agent’s payoff is 2, it is more significantly rewarded when their value is 3 as opposed to when their value is 9. This design ensures that the utility function remains concave for all agents.
@@ 3.2 Q-Learning
where $Q(s,a)$ is the Q-value function that estimates the expected future reward for taking action $a$ in state $s$, $t$ is the current time step, and $\arg\max_{a}Q(s,a)$ is the action that maximizes the Q-value function for the current state $s$. The exploration rate at time $t$ is determined by $\epsilon_{t}=\epsilon_{\text{max}}\cdot\text{decay\_rate}^{t}$, where $\epsilon_{\text{max}}$ is the maximum exploration rate (i.e., the initial value of $\epsilon$), decay_rate is a constant $[0,1]$ that determines the rate of decay, and $t$ is the current time step.
While the scenarios presented in this paper (i.e., auctions with a single bidding round and one-step episodes) differ from traditional RL problems (with numerous steps per episode), the use of Q-Learning and the Bellman equation remains valid. Even though the traditional interpretation of the Bellman equation involves considering future rewards and actions over multiple steps, in a single-step scenario, the equation simplifies to focus on the immediate reward and action taken.
@@ 3.3.1 Deep Policy Gradient
The DNN used in Deep Policy Gradient approaches takes the current state of the environment as an input and outputs a probability distribution over actions (for discrete action spaces) or parameters of a distribution from which actions are sampled (for continuous action spaces). The DNN’s weights and biases are the policy parameters that are then iteratively adjusted during training. Sophisticated optimisers can also be used with Deep Policy methods, which handle noisy gradient estimates much better than VPG approaches. An entropy term is commonly used with the loss function to assist convergence and avoid sub-optimal policies.
@@ 3.4 Actor-Critic
The Actor-Critic approach combines aspects of both policy-based (the “Actor”) and value-based (the “Critic”) learning. The actor component is responsible for selecting actions based on the current policy, that is, a mapping from states to actions. The critic evaluates the actions taken by the actor by estimating the value function, which measures the expected return from a given state using the current policy. The critic’s evaluation is then used to update the actor’s policy towards more rewarding actions. One key advantage of actor-critic methods is their ability to handle continuous and discrete action spaces. In addition to this versatility, separating the policy and value function estimates can stabilize the learning process when compared to methods using only policy or value function estimation.
@@ 3.4.1 Advantage Actor-Critic
Mathematically, the advantage function, $A(s,a)$, is defined as the difference between the action-value function, $Q(s,a)$, and the state-value function, $V(s)$:
@@
A2C excels in scenarios where future rewards must be estimated, and decisions must consider long-term outcomes. Consequently, the advantage function plays a crucial role in A2C methods, facilitating the balance between immediate and future rewards. In the single-bid auction scenarios examined in this paper, such features are underutilised and possibly unnecessary. Each auction episode is effectively a single decision with an immediate reward, and the learning process primarily focuses on associating states directly with their resulting rewards without considering future states. Nonetheless, experimental results using A2C (see section 4) reveal a substantial improvement over simpler policy gradient and Q-learning methods, especially regarding learning stability.
Table 1: Main advantages and disadvantages of the six reinforcement learning algorithms considered in the paper.
@@ 3.4.2 Proximal Policy Optimisation
Proximal Policy Optimization (PPO) was introduced in 2017 by John Schulman and colleagues in their paper “Proximal Policy Optimization Algorithms” (Schulman et al., 2017). This paper presented PPO as a simpler alternative to Trust Region Policy Optimization (TRPO) while achieving similar or better performance. PPO works by limiting the change in policy during each update to avoid disruptively large policy updates.
In the context of reinforcement learning (RL) and specifically Proximal Policy Optimization (PPO), a policy refers to an agent’s strategy to make decisions. It defines the mapping from perceived states of the environment to actions to be taken. In the case of PPO, the algorithm trains a stochastic policy in an on-policy way, which means that it explores by sampling actions according to the latest version of its stochastic policy
The policy in RL can be deterministic or stochastic. A deterministic policy directly maps states to actions, while a stochastic policy outputs a probability distribution over actions for a given state. PPO focuses on training a stochastic policy, enhancing exploration and robustness when applied to complex environments. The PPO algorithm’s objective is to improve the policy in a stable and sample-efficient manner, enabling the agent to make better decisions over time.
@@
The key idea behind PPO is to improve the training stability of the policy by limiting the change made to the policy during each update. This is achieved by constraining the policy updates to a ‘trust region’, which prevents the new policy from deviating too far from the old policy. In an auction environment with private information and non-deterministic outcomes, minimising the size of policy updates is an important feature. Unpredictable by other auction participants (agent or human) during the training of an RL agent can severely compromise learning by misleading an algorithm’s search for the best decision. PPO achieves this via a clipped surrogate objective function. The objective of this function is defined as:
@@
Like A2C, using PPO in single-bid auctions does not fully exploit the algorithm’s potential. PPO is designed for environments where actions have long-term consequences over multiple steps; however, as displayed in section 4 below, PPO performs exceptionally well in all auction types examined in this paper.
In summary, Table 1 provides an account of the advantages and disadvantages of each RL algorithm used in this paper. The following section then presents the results from a series of empirical experiments with each of the six algorithms performing as bidders in the three different auction types: Discretionary Price, Generalised Second-Price, and Uniform-Price.
@@ 4 Simulations
In this section, we conduct various simulations using the six algorithms described above. Since the primary aim of this paper is to assess various RL algorithms that are suitable for simulating bidding behavior, we maintain a consistent number of bidders across all simulations and only vary the total number of units available to allow for a tractable analysis. We assume each auction features six bidders, each demanding two units. We explore three scenarios with regard to the total units available for each auction: one with four total units, one with six units, and one with eight units. [^2] These numbers are chosen to represent three cases: one where the total number of units available is below half of the total demand, one where it is equal to half, and one where it is more than half but less than the total demand (Engelbrecht-Wiggans and Kahn, 1998; Khezr and Menezes, 2017).
The main variable used to assess the learning process in each simulation is called the ‘learning ratio’, which is calculated as the value minus the bid, divided by the value. Intuitively, one expects an agent following an optimal strategy to submit bids that proportionally differ from their value in a stable manner. Therefore, the learning ratio is a good metric to test such performance. Once we confirm that the learning patterns demonstrate reasonably consistent learning, we then use revenue and efficiency as the metrics to evaluate auction performance. Revenue is simply the sum of payments made by all winning bidders, and efficiency is calculated as the sum of the $K$ highest values (irrespective of the bids), where $K$ is the number of units available minus the sum of the values of those who won an item. When this term is zero, the auction has allocated the objects in a fully efficient manner.

### v3

@@ Abstract
Understanding bidding behavior in multi-unit auctions remains an ongoing challenge for researchers. Despite their widespread use, theoretical insights into the bidding behavior, revenue ranking, and efficiency of commonly used multi-unit auctions are limited. This paper utilizes artificial intelligence, specifically model-free reinforcement learning, to simulate bidding in three prominent multi-unit auctions employed in practice. We introduce six algorithms that are suitable for learning and bidding in multi-unit auctions and compare them using an illustrative example. This paper underscores the significance of using artificial intelligence in auction design, particularly in enhancing the design of multi-unit auctions.
@@ 1 Introduction
Multi-unit auctions play a fundamental role in the allocation of goods and services across various markets, including treasury bills, emission permits, spectrum licenses, and electricity (Khezr and Cumpston, 2022; Bichler and Goeree, 2017). These auctions are often characterized by asymmetric information and, due to the existence of multiple units, result in a complex bidding environment. Although multi-unit auctions are commonly used in practice, there is a lack of consensus in the literature regarding their performance in terms of revenue and efficiency. Moreover, economic theory often lacks clear guidance on the outcomes of many multi-unit auctions (Krishna, 2009).
A range of disciplines, including economics and computer science, have applied various methodologies to investigate multi-unit auctions. Auction theory, data analytics, experimental approaches, and simulations represent some of the predominant methods used in these studies (Ausubel et al., 2014; Roughgarden et al., 2017; Morgenstern and Roughgarden, 2016; Hailu and Thoyer, 2007). Each method offers distinct advantages but also comes with limitations. A common flaw in these approaches is their inability to comparatively assess the performance of different auction variations under similar conditions. Additionally, traditional simulation methods often lack clear guidance as they require an existing model or theoretical prediction to feed into the simulation. In this article, we use artificial intelligence, specifically introducing several reinforcement learning (RL) algorithms, to simulate bidding behavior in various multi-unit auctions.
@@
Our approach leverages artificial intelligence as a complementary tool to previous methodologies, aiming to emulate the human thought process and decision-making in complex bidding scenarios within multi-unit auctions. Using a standard private value model, we assume that buyers’ valuations for each unit up for auction are privately known. Our investigation focuses on three sealed-bid multi-unit auctions: the Discriminatory Price (DP), the Generalized Second-Price (GSP), and the Uniform-Price (UP) auctions. All three of these auctions have been extensively used in real-world markets to sell various goods and services.
We first provide a metric for learning and demonstrate through an example, in which each bidder demands two units, that almost all RL algorithms perform very well in terms of learning to bid. Although they vary in terms of speed and learning patterns, some top-performing algorithms surprisingly converge to very similar strategies in the final episodes of learning. One of the six algorithms, namely Proximal Policy Optimization, not only performs the best by earning the highest payoffs in all cases, but is also the most stable in terms of learning. We then compare the revenue and efficiency of the three auctions in three different scenarios.
@@
Since the seminal work by Vickrey (Vickrey, 1961), it is known that bidding behavior in multi-unit auctions does not necessarily follow that in single-unit auctions. Despite this, most conventionally used multi-unit auctions are extensions of a standard single-unit format. For instance, the generalized second-price auction, as appears from its name, is a format that extends the second-price rule to a multi-unit case. There are several studies that show these multi-unit extensions do not necessarily have the ideal properties of the single unit format (e.g. Back and Zender, 1993; Edelman et al., 2007; Ausubel et al., 2014). Additionally, due to multiple equilibria and lack of closed-form solutions for the bidding functions, theoretical studies have no clear conclusion about the revenue and efficiency ranking of commonly used multi-unit auctions (Krishna, 2009; Ausubel et al., 2014).
Recently, the use of RL approaches, particularly Q-learning algorithms, has become more prevalent in the auction theory literature. For example, Banchio and Skrzypacz (2022) employ a simple Q-learning algorithm to simulate bidding behavior in two different single-unit auction formats. Khezr et al. (2024) is another study that uses Q-learning to simulate a more complex type of auction, namely the knapsack auction. While various other RL approaches could potentially be used to simulate bidding in auctions, current studies are mainly limited to simple Q-learning algorithms. Therefore, our main aim is to explore other, more advanced algorithms that could effectively simulate such complex bidding environments.
@@ 2 Model
The seller uses a sealed-bid multi-unit auction to allocate the objects. Each bidder submits a vector of bids $\mathbf{b}^{i}$ for all units they demand. The auctioneer ranks the bids from the highest to the lowest and allocates the object to the $K$ highest submitted bids. We explore three types of payment rules as follows.
@@
In the discriminatory price auction, also known as pay as bid auction, every buyer $i$ submits up to $k$ bids, represented by $\mathbf{b}^{1}=(b_{1}^{i},b_{2}^{i},...,b_{k}^{i})$. The auctioneer arranges all the $n\times k$ bids from the highest to the lowest. The top $k$ bids each win a unit and pay the amount they bid for that particular unit. For example, if the first bid from bidder three, $b_{1}^{3}$, holds the top rank, she secures the first unit and pays $b_{1}^{3}$ for it.
@@
Similarly to the DP auction, each buyer $i$ submits up to $k$ bids. The top $k$ bids each win a unit. However, in this case, winners pay the amount of the bid that is ranked immediately below their own (excluding their own subsequent bids). For instance, if bidder three’s first bid, $b_{1}^{3}$, holds the highest rank, she secures the first unit but pays the amount of the second-highest bid.
@@
The Uniform Price Auction, as one of the most popular auctions in practice, aims to replicate competitive equilibrium pricing by setting a uniform price for all units won in the auction. Similar to the previous two auctions, the top $k$ bids each win a unit. However, in this auction, every winning bidder pays a uniform price for each unit won, which is equal to the highest losing bid—that is, the $K+1$th highest bid.
@@ 3 Reinforcement Learning approaches
We use Reinforcement Learning (RL) methods to model bidder behavior in multi-unit auctions with multiple bidders. Inspired by behavioral psychology, RL algorithms guide an agent to learn how to accomplish a goal via feedback (consisting of rewards and penalties). An agent, in the context of this work, is a simulated human bidder operating within an environment, an auction, in which they interact with other agents via bidding for available items. The environment encompasses everything the agent interacts with and can change over time, either due to the agent’s actions or independently. Interactions between an agent and the environment are defined by the terms state, action, and reward.
A state is a complete description of the agent’s situation within the environment. An observation, which might be a partial view of the state, provides the agent with information about the current situation. Actions are the set of all possible decisions the agent can make. A reward is a feedback signal indicating the direct effect of the agent’s previous action. The agent’s objective is to maximize their return (or cumulative reward), over time. This objective and the associated decision-making strategy are encapsulated in an agent’s policy, a mapping of states (or observations) to actions.
@@ 3.1 Environment setup and evaluation
The reward function, therefore, returns values that can be either positive or negative, rewarding positive payoffs and punishing overbidding more severely than unsuccessful bidding. Specifically, when $p>0$, the agent receives a positive reward as determined by Equation 1. However, when $p<0$, it indicates that the agent’s payment—whether it is their own bid (in DP) or some other lower bid (in GSP and UP)—exceeds their value thus the numerator of the reward function in Equation 1 becomes negative. In this case, the agent receives a penalty in the form of a negative reward. Both reward and punishment are proportional to the value, which means that for higher values, any given positive or negative payoff is less rewarded or punished respectively. For instance, if an agent’s payoff is 2, it is more significantly rewarded when their value is 3 as opposed to when their value is 9. This design ensures that the utility function remains concave for all agents.
@@ 3.2 Q-Learning
where $Q(s,a)$ is the Q-value function that estimates the expected future reward for taking action $a$ in state $s$, $t$ is the current time step, and $\arg\max_{a}Q(s,a)$ represents the action that maximises the Q-value function in the current state $s$. The exploration rate at time $t$ is determined by $\epsilon_{t}=\epsilon_{\text{max}}\cdot\text{decay\_rate}^{t}$, where $\epsilon_{\text{max}}$ is the maximum exploration rate (i.e., the initial value of $\epsilon$), decay_rate is a constant $[0,1]$ that determines the rate of decay, and $t$ is the current time step.
While the scenarios presented in this paper (i.e., auctions with a single bidding round and one-step episodes) differ from traditional RL problems (with numerous steps per episode), the use of Q-Learning and the Bellman equation remains valid. Although the traditional interpretation of the Bellman equation involves considering future rewards and actions over multiple steps, in a single-step scenario, the equation simplifies to focus on the immediate reward and action taken.
@@ 3.3.1 Deep Policy Gradient
Deep Policy Gradient approaches employ a DNN using the current environment state as an input, and produces either a probability distribution over actions as an output (for discrete action spaces), or parameters of a distribution from which actions are sampled (for continuous action spaces). The DNN’s weights and biases are the policy parameters that are then iteratively adjusted during training. Sophisticated optimisers can also be used with Deep Policy methods, which handle noisy gradient estimates much better than VPG approaches. An entropy term is commonly used with the loss function to assist convergence and avoid sub-optimal policies.
@@ 3.4 Actor-Critic
The Actor-Critic approach combines aspects of both policy-based (the “Actor”) and value-based (the “Critic”) learning. The actor component is responsible for selecting actions based on the current policy, that is, a mapping from states to actions. The role of the Critic is to assess the actions taken by the Actor through value function estimation that measures the expected return from a given state using the current policy. The critic’s evaluation is then used to update the actor’s policy towards more rewarding actions. One key advantage of actor-critic methods is their ability to handle continuous and discrete action spaces. In addition to this versatility, separating the policy and value function estimates can stabilize the learning process when compared to methods using only policy or value function estimation.
@@ 3.4.1 Advantage Actor-Critic
An advantage function, $A(s,a)$, encapsulates the difference between the action-value function, $Q(s,a)$, and the state-value function, $V(s)$:
@@
A2C excels in scenarios where future rewards must be estimated, and decisions must consider long-term outcomes. Consequently, the advantage function plays a crucial role in A2C methods, facilitating the balance between immediate and future rewards. In the single-bid auction scenarios examined in this paper, such features are underutilised and possibly unnecessary. Each auction episode is effectively a single decision with an immediate reward, and the learning process focuses primarily on associating states directly with their resulting rewards without considering future states. Nonetheless, experimental results using A2C (see section 4) reveal a substantial improvement over simpler policy gradient and Q-learning methods, especially regarding learning stability.
Table 1: Principal advantages and disadvantages of the six reinforcement learning algorithms considered in the paper.
@@ 3.4.2 Proximal Policy Optimisation
Proximal Policy Optimization (PPO) was introduced in 2017 by John Schulman and colleagues in their paper ’Proximal Policy Optimization Algorithms’ (Schulman et al., 2017). This paper presented PPO as a simpler alternative to Trust Region Policy Optimization (TRPO) while achieving similar or better performance. PPO works by limiting the change in policy during each update to avoid disruptively large policy updates.
In both RL in general and PPO in particular, a policy refers to an agent’s strategy for making decisions. It defines the mapping from perceived environmental states to actions to be taken. In the case of PPO, the algorithm trains a stochastic policy using an on-policy method, whereby exploration is achieved by sampling actions according to the latest version of its stochastic policy.
The policy in RL can be deterministic or stochastic. A deterministic policy directly maps states to actions, while a stochastic policy outputs a probability distribution over actions for a given state. PPO focuses on training a stochastic policy, enhancing exploration and robustness when applied to complex environments. The objective of the PPO algorithm is to improve the policy in a stable and sample-efficient manner, enabling the agent to make better decisions over time.
@@
The key idea behind PPO is to limit the degree of policy change during updates to improve training stability. This is achieved by constraining the policy updates to a ‘trust region’, which prevents the new policy from deviating too far from the old policy. In an auction environment with private information and non-deterministic outcomes, minimizing the size of policy updates is an important feature. Unpredictable by other auction participants (agent or human) during the training of an RL agent can severely compromise learning by misleading an algorithm’s search for the best decision. PPO achieves this via a clipped surrogate objective function. The objective of this function is defined as:
@@
Like A2C, using PPO in single-bid auctions does not fully exploit the algorithm’s potential. PPO is designed for environments where actions have long-term consequences over multiple steps; however, as shown in Section 4 below, PPO performs exceptionally well in all auction types examined in this paper.
In summary, Table 1 provides an account of the advantages and disadvantages of each RL algorithm used in this paper. The following section then presents the results from a series of empirical experiments with each of the six algorithms performing as bidders in the three different auction types: Discretionary price, generalized second price, and uniform price.
@@ 4 Simulations
In this section, we conduct various simulations using the six algorithms described above. Since the primary aim of this paper is to assess various RL algorithms that are suitable for simulating bidding behavior, we maintain a consistent number of bidders across all simulations and only vary the total number of units available to allow for a tractable analysis. We assume that each auction features six bidders, each demanding two units. We explore three scenarios with respect to the total units available for each auction: one with four total units, one with six units, and one with eight units. [^2] These numbers are chosen to guarantees that we account for all possible equilibria (Engelbrecht-Wiggans and Kahn, 1998; Khezr and Menezes, 2017). The values of each agent for both units are the same and are random draws from a uniform distribution on $[0,10]$.
The main variable used to assess the learning process in each simulation is called the ‘learning ratio’, which is calculated as the value minus the bid divided by the value. Intuitively, one expects an agent following an optimal strategy to submit bids that proportionally differ from their value in a stable manner. Therefore, the learning ratio is a good metric to test this performance. Once we confirm that the learning patterns demonstrate reasonably consistent learning, we then use revenue and efficiency as the metrics to evaluate auction performance. Revenue is simply the sum of payments made by all winning bidders, and efficiency is calculated as the sum of the $K$ highest values (irrespective of the bids), where $K$ is the number of units available minus the sum of the values of those who won an item. When this term is zero, the auction has allocated the objects in a fully efficient manner.